# On the Robustness of Kolmogorov-Arnold Networks: An Adversarial Perspective

### v1

@@ 3.4 Analysis of White-Box Attacks
Table 3: White-box attacks on KANs and MLPs across MNIST, FashionMNIST, and KMNIST datasets. The attacks are performed using the $\ell_{\infty}$ norm with $\epsilon=32/255$ for all datasets. Robust accuracy is given for each corresponding attack. The “Acc” column refers to the clean accuracy of the model. Boldface values indicate the most robust model in each experiment.
Dataset Model FGSM PGD C&W MIM Acc
MNIST $\text{KAN}_{\text{small}}$ 6.9 0.0 0.0 0.0 96.0
$\text{MLP}_{\text{small}}$ 1.4 0.0 0.5 0.6 97.9
$\text{KAN}_{\text{medium}}$ 9.5 0.0 0.0 0.0 97.6
$\text{MLP}_{\text{medium}}$ 10.8 0.8 0.9 1.3 98.3
$\text{KAN}_{\text{large}}$ 37.8 3.2 4.1 0.9 94.5
$\text{MLP}_{\text{large}}$ 9.6 4.9 0.2 4.9 97.5
FashionMNIST $\text{KAN}_{\text{small}}$ 4.8 0.0 0.0 0.0 86.9
$\text{MLP}_{\text{small}}$ 0.9 0.2 0.4 0.4 88.2
$\text{KAN}_{\text{medium}}$ 3.8 0.0 0.0 0.0 86.7
$\text{MLP}_{\text{medium}}$ 5.2 0.2 0.1 0.3 89.2
$\text{KAN}_{\text{large}}$ 16.7 0.2 0.3 0.0 83.2
$\text{MLP}_{\text{large}}$ 1.5 0.0 0.1 0.1 88.5
KMNIST $\text{KAN}_{\text{small}}$ 2.8 0.0 0.0 0.0 81.7
$\text{MLP}_{\text{small}}$ 3.1 1.2 1.5 1.5 90.3
$\text{KAN}_{\text{medium}}$ 7.6 0.0 0.0 0.0 87.5
$\text{MLP}_{\text{medium}}$ 11.7 1.0 1.1 1.6 91.1
$\text{KAN}_{\text{large}}$ 25.9 0.9 1.2 0.1 78.9
$\text{MLP}_{\text{large}}$ 7.1 0.8 0.3 1.2 89.3
In the context of the iterative adversarial attack methods (PGD, C&W, MIM), small and medium KAN models exhibited lower robustness, when compared with small and medium MLP models as shown in Table 3. Across all examined datasets, medium KAN model displayed a lack of resistance against the deployed attacks. A similar trend was observed in small KAN models, which generally showed almost no robustness, except in the specific case of the MNIST dataset.
In contrast, small and medium MLPs exhibited a certain degree of resistance against these attacks. The sole exception to this trend was observed for the small-sized MLPs with the MNIST dataset, which were completely vulnerable, failing to correctly classify any adversarial images.
@@
The trend whereby MLPs generally exhibit greater robustness compared to KANs changes when considering the larger configurations. For large models, the robustness of KAN improves, surpassing that of large MLP under the C&W attack across all the datasets, and under the PGD attack on FashionMNIST and KMNIST datasets. Not only do large KANs surpass large MLPs in robustness under some iterative-adversarial attacks, but they also exhibited marked improvement when compared to smaller and medium KANs. Across all iterative adversarial attacks and datasets, large KAN showed significantly greater resilience than both small and medium KANs.
@@
Further evidence supporting the robustness of the large KAN model can be observed in Figure 1. The figure illustrates that the loss function (which increases as the discrepancy between the predicted label and the true label is maximized), is consistently lower for the large KAN compared to the other models. This indicates that it is more difficult to ‘deceive’ this model. Additionally, another indicator of this model’s resilience is that, while the losses of other models tend to converge, the loss of the large KAN remains linear.
@@
Figure 1: Comparative loss dynamics of MLPs and KANs under PGD, MIM, and C&W attacks on the MNIST, FashionMNIST, and KMNIST datasets (from top to bottom, respectively). Each line represents the mean loss across batches per iteration, with shaded areas indicating the standard deviation.
In the context of the non-iterative attack method (FGSM), results were somewhat mixed when comparing small and medium models; however, a distinct pattern emerged for larger models, as seen in Table 3: large KAN significantly outperformed large MLP, demonstrating greater robustness to FGSM attacks. To further illustrate the differneces between the medium and large MLPs and KANs, we conducted additional tests using the FGSM attack, focusing on how robustness changes as the $\epsilon$ value increases.
The results, shown in Figure 2, clearly demonstrate that at small $\epsilon$ values, the robustness of both MLPs and KANs appears relatively similar. However, as epsilon increases, the large KAN model is the only one that maintains a significant level of robustness, distinguishing it from medium KAN and both MLP models. Note that at very high $\epsilon$ values medium KAN exhibits greater robustness than even large MLP.
@@
Figure 2: Robust accuracy of $\text{MLP}_{\text{medium}}$, $\text{MLP}_{\text{large}}$, $\text{KAN}_{\text{medium}}$, and $\text{MLP}_{\text{large}}$ as a function of FGSM attack strength, with varying $\epsilon$ values. Each bar represents a different model, and robustness is measured as the accuracy of the model against adversarial examples generated with specific $\epsilon$ values. The x-axis indicates the $\epsilon$ values used for FGSM attacks, while the y-axis shows the corresponding robust accuracy of the models.
Despite the observed improvements in robustness with increased model size, KANs remain notably fragile to adversarial attacks. The results presented in Table 3 and Figure 1 indicate that even the large KAN models, while more resilient than their smaller counterparts and sometimes outperforming large MLPs, still exhibit significant vulnerability under adversarial conditions. This fragility is particularly pronounced in smaller and medium KAN configurations, which show almost no resistance against certain types of attacks, such as the PGD and MIM methods.
Appendix A shows sample image attacks.
4 Transferability Study
How transferable are the adversarial examples created by KANs?
As demonstrated in Section 3, white-box attacks are highly effective at generating adversarial examples that successfully deceive KANs. We now extend the previous analysis by examining the transferability of adversarial examples that are misclassified by KANs. Transferability refers to the phenomenon whereby adversarial examples are misclassified by multiple classifiers.
The transferability of adversarial examples has been extensively documented for various deep-learning architectures. The concept of adversarial-example transferability was first introduced by Szegedy et al. [41]. Subsequent studies have explored the transferability of adversarial examples across different architectures [47, 46, 51, 49, 1, 10, 32]. In this section, we provide a detailed evaluation and analysis of this phenomenon.
4.1 Transferability Metric

### v2

@@ 4.1 Fully Connected Models
Table 4: White-box and black-box attacks on fully connected models evaluated across MNIST, FashionMNIST, and KMNIST datasets. Robust accuracy is given for each corresponding attack. The “Acc” column refers to the clean accuracy of the model. Boldface values indicate the most robust model in each experiment.
Dataset Model White-Box Black-Box Acc
FGSM PGD C&W MIM Square SimBA NES
MNIST $\text{FCKAN}_{\text{small}}$ 6.96 0.00 0.01 0.00 0.09 0.00 0.00 96.02
$\text{FCNN}_{\text{small}}$ 1.40 0.52 0.65 0.62 2.38 0.19 2.08 97.94
$\text{FCKAN}_{\text{medium}}$ 9.51 0.00 0.00 0.00 0.00 0.19 0.00 97.62
$\text{FCNN}_{\text{medium}}$ 10.80 0.77 0.87 1.05 3.27 18.35 19.94 98.31
$\text{FCKAN}_{\text{large}}$ 37.83 1.64 2.27 0.42 24.20 38.78 0.00 94.54
$\text{FCNN}_{\text{large}}$ 9.57 4.78 0.15 4.76 2.38 5.75 5.25 97.47
FashionMNIST $\text{FCKAN}_{\text{small}}$ 4.84 0.00 0.00 0.00 0.00 0.00 0.00 86.83
$\text{FCNN}_{\text{small}}$ 0.95 0.23 0.44 0.32 3.86 0.39 3.67 88.20
$\text{FCKAN}_{\text{medium}}$ 3.78 0.00 0.00 0.00 0.00 0.00 0.00 86.72
$\text{FCNN}_{\text{medium}}$ 5.23 0.12 0.11 0.18 1.58 11.30 11.80 89.14
$\text{FCKAN}_{\text{large}}$ 16.68 0.05 0.09 0.00 19.64 29.06 0.00 83.24
$\text{FCNN}_{\text{large}}$ 1.48 0.11 0.00 0.11 0.19 2.48 2.67 88.53
KMNIST $\text{FCKAN}_{\text{small}}$ 2.80 0.00 0.00 0.00 0.00 0.00 0.00 81.71
$\text{FCNN}_{\text{small}}$ 3.09 1.16 1.52 1.41 7.24 0.89 5.55 90.28
$\text{FCKAN}_{\text{medium}}$ 7.65 0.00 0.00 0.00 0.00 1.48 1.68 87.47
$\text{FCNN}_{\text{medium}}$ 11.74 1.01 1.11 1.35 6.54 8.82 17.85 91.15
$\text{FCKAN}_{\text{large}}$ 25.90 0.36 0.53 0.04 14.78 23.61 0.00 78.92
$\text{FCNN}_{\text{large}}$ 7.14 0.70 0.29 1.05 3.86 1.38 2.67 89.28
@@
This trend, in which small and medium FCNNs generally outperform FCKANs in robustness under iterative white-box attacks, and for which FCKANs often failed to classify any adversarial images generated by these methods (except for the small FCKAN on MNIST), begins to shift when considering larger configurations. For large models, the robustness of FCKANs improves significantly, surpassing that of large FCNNs under the C&W attack across all datasets. Not only do large FCKANs outperform large FCNNs under the C&W attack, but they also show substantial improvement over their smaller and medium-sized counterparts. Across all iterative adversarial attacks and datasets, large FCKANs exhibited considerably greater resilience than small and medium FCKANs. In contrast, large FCNNs often demonstrated lower robustness than their smaller variants.
@@
Further evidence of the robustness of large FCKANs under white-box iterative attacks is shown in Figure 1. The figure illustrates that the loss function, which increases as the discrepancy between the predicted and true labels is maximized, remains consistently lower for large FCKANs compared to other models. This indicates that deceiving large FCKANs is more challenging. Additionally, while the losses of other models tend to converge during attacks, the loss of large FCKANs maintains a linear trajectory, further underscoring their resilience.
@@
Figure 1: Comparative loss dynamics of FCNNs and FCKANs under PGD, MIM, and C&W attacks on the MNIST, FashionMNIST, and KMNIST datasets (from top to bottom, respectively). Each line represents the mean loss across batches per iteration, with shaded areas indicating the standard deviation.
A particularly notable observation is the superior performance of large FCKANs compared to large FCNNs under the FGSM white-box adversarial attack, with a pronounced gap favoring FCKANs across all datasets. This is further supported by Figure 2, which illustrates the clear advantage of large FCKANs. While the differences in robustness between small and medium FCNNs and FCKANs are less definitive, the superiority of large FCKANs over large FCNNs is both clear and substantial, firmly establishing FCKANs as the more robust architecture in larger configurations.
@@
Figure 2: Robust accuracy of $\text{FCNN}_{\text{medium}}$, $\text{FCNN}_{\text{large}}$, $\text{FCKAN}_{\text{medium}}$, and $\text{FCKAN}_{\text{large}}$ as a function of FGSM attack strength, with varying $\epsilon$ values. Each bar represents a different model, and robustness is measured as the accuracy of the model against adversarial examples generated with specific $\epsilon$ values. The x-axis indicates the $\epsilon$ values used for FGSM attacks, while the y-axis shows the corresponding robust accuracy of the models.
When considering black-box attacks (Square, SimBA, and NES), the performance dynamics shift. Small FCNNs outperform FCKANs on all datasets, where small FCKANs failed to classify any image generated with those attacks. When considering the medium models, although FCKANs successfully classified a small number of images (especially on the KMNIST dataset), medium FCNNs still performed better than medium FCKANs across all datasets. However, for larger models, FCKANs demonstrate a marked improvement in robustness against SimBA and Square black-box attacks, significantly surpassing FCNNs. In contrast, similar to the small and medium sizes under NES attack, FCKANs failed to classify any adversarial image. The fact that FCKANs failed under the NES attack could result from the gradient approximation for these architectures being much more accurate than for the FCNN architecture.
In our analysis of the transferability of adversarial white-box attacks (Table 5) , we observed that across all datasets, the large FCKAN consistently exhibited the lowest percentage of misclassified images among all six models. This suggests that fooling the large FCKAN is the most challenging task. Moreover, except for the small and medium models on the KMNIST dataset, FCKANs consistently demonstrated greater resistance to adversarial attacks than their FCNN counterparts for each model size (small, medium and large), highlighting the superior robustness of FCKANs at every scale.
Table 5: Maximum transferability of our fully connected models, for FashionMNIST, KMNIST, and MNIST. The value in row $i$, column $j$ represents the maximum transferability between row-$i$ model and column-$j$ model across four attacks—FGSM, PGD, C&W, MIM—calculated using Equation 11. Unlike robustness tables, the values here indicate the percentage of images misclassified. The ‘Average’ row shows the average transferability for the attacking model. Boldfaced values denote the lowest transferability, where a lower value indicates better robustness.
(a) FashionMNIST
$\text{FCKAN}_{\text{small}}$ $\text{FCKAN}_{\text{medium}}$ $\text{FCKAN}_{\text{large}}$ $\text{FCNN}_{\text{small}}$ $\text{FCNN}_{\text{medium}}$ $\text{FCNN}_{\text{large}}$
$\text{FCKAN}_{\text{small}}$ 100.00 57.32 38.97 59.98 56.23 48.76
$\text{FCKAN}_{\text{medium}}$ 40.55 100.00 28.11 39.89 39.99 35.81
$\text{FCKAN}_{\text{large}}$ 32.34 29.91 100.00 31.90 26.03 27.18
$\text{FCNN}_{\text{small}}$ 81.34 81.57 59.43 99.76 98.46 90.64
$\text{FCNN}_{\text{medium}}$ 64.82 68.91 45.16 90.19 99.86 85.33
$\text{FCNN}_{\text{large}}$ 57.13 59.31 39.45 75.18 82.21 100.00
Average ($\downarrow$) 62.69 66.16 51.85 66.15 67.13 64.61
(b) KMNIST
$\text{FCKAN}_{\text{small}}$ $\text{FCKAN}_{\text{medium}}$ $\text{FCKAN}_{\text{large}}$ $\text{FCNN}_{\text{small}}$ $\text{FCNN}_{\text{medium}}$ $\text{FCNN}_{\text{large}}$
$\text{FCKAN}_{\text{small}}$ 100.00 29.58 21.11 21.60 13.77 19.71
$\text{FCKAN}_{\text{medium}}$ 32.87 100.00 17.47 16.19 10.94 17.43
$\text{FCKAN}_{\text{large}}$ 19.59 11.82 99.94 7.99 4.69 11.30
$\text{FCNN}_{\text{small}}$ 70.68 62.29 45.95 98.72 89.77 71.58
$\text{FCNN}_{\text{medium}}$ 51.04 42.42 30.76 83.49 98.88 59.40
$\text{FCNN}_{\text{large}}$ 45.75 35.72 31.89 50.22 43.74 99.66
Average ($\downarrow$) 53.32 46.97 41.18 46.36 43.46 46.51
(c) MNIST
$\text{FCKAN}_{\text{small}}$ $\text{FCKAN}_{\text{medium}}$ $\text{FCKAN}_{\text{large}}$ $\text{FCNN}_{\text{small}}$ $\text{FCNN}_{\text{medium}}$ $\text{FCNN}_{\text{large}}$
$\text{FCKAN}_{\text{small}}$ 100.00 28.10 22.94 55.10 30.55 32.32
$\text{FCKAN}_{\text{medium}}$ 35.33 100.00 17.71 38.83 22.67 26.74
$\text{FCKAN}_{\text{large}}$ 22.78 08.94 99.76 18.50 08.36 14.56
$\text{FCNN}_{\text{small}}$ 81.35 68.48 48.68 99.78 95.84 88.31
$\text{FCNN}_{\text{medium}}$ 64.17 54.40 33.83 95.48 99.62 84.39
$\text{FCNN}_{\text{large}}$ 58.89 44.78 31.77 78.42 72.29 99.92
Average ($\downarrow$) 60.41 50.78 42.44 64.35 54.88 57.70
Qualitative results corresponding to these attacks are presented in Appendix A.
4.2 Convolutional Models
In the context of iterative white-box adversarial attacks, CKANs and CNNs exhibit contrasting behaviors across datasets. As observed in Table 6, both CKANs and CNNs exhibit poor performance on the complex datases (CIFAR-10 and SVHN), demonstrating significant vulnerability to adversarial perturbations. On CIFAR-10, none of the models successfully classify any images generated by those attacks. On the SVHN dataset, the medium and large models manage to classify a small number of adversarially generated images. This suggests that as the dataset complexity increases, the models become more susceptible to adversarial attacks, likely due to the intricate decision boundaries required for accurate classification.
Table 6: White-box and black-box attacks on convolutional models, evaluated across MNIST, CIFAR-10, and SVHN datasets. The table reports robust accuracy for each attack, while the ‘Acc‘ column indicates the model’s accuracy on clean, unperturbed inputs. Boldface values indicate the most robust model in each experiment.
Dataset Model White-Box Black-Box Acc
FGSM PGD C&W MIM Square SimBA NES
MNIST $\text{CKAN}_{\text{small}}$ 72.36 4.52 5.70 8.98 38.39 34.82 49.90 99.47
$\text{CNN}_{\text{small}}$ 68.20 0.74 0.98 2.26 27.38 9.82 33.73 99.23
$\text{CKAN}_{\text{medium}}$ 86.33 29.88 33.12 42.72 57.93 41.76 66.07 99.55
$\text{CNN}_{\text{medium}}$ 28.36 0.00 0.00 0.00 0.39 7.24 8.92 99.36
$\text{CKAN}_{\text{large}}$ 84.72 31.28 34.32 39.64 56.84 36.21 58.92 99.52
$\text{CNN}_{\text{large}}$ 55.87 0.01 0.13 0.19 5.15 2.38 5.45 99.49
CIFAR-10 $\text{CKAN}_{\text{small}}$ 0.01 0.00 0.00 0.00 0.00 2.48 0.99 74.70
$\text{CNN}_{\text{small}}$ 0.01 0.00 0.00 0.00 0.00 2.28 0.69 73.03
$\text{CKAN}_{\text{medium}}$ 0.57 0.00 0.00 0.00 0.39 4.26 1.88 80.72
$\text{CNN}_{\text{medium}}$ 0.27 0.00 0.00 0.00 0.09 2.28 0.79 76.43
$\text{CKAN}_{\text{large}}$ 0.85 0.00 0.00 0.00 0.19 4.06 1.98 82.82
$\text{CNN}_{\text{large}}$ 1.94 0.00 0.00 0.00 1.48 4.46 2.38 79.26
SVHN $\text{CKAN}_{\text{small}}$ 1.04 0.00 0.00 0.00 7.44 10.41 7.44 92.34
$\text{CNN}_{\text{small}}$ 0.43 0.00 0.00 0.00 4.76 5.15 2.87 90.69
$\text{CKAN}_{\text{medium}}$ 6.74 0.02 0.02 0.02 16.96 16.36 13.59 94.37
$\text{CNN}_{\text{medium}}$ 5.00 0.05 0.05 0.04 14.78 9.52 8.03 93.41
$\text{CKAN}_{\text{large}}$ 8.33 0.09 0.10 0.09 19.84 16.26 13.09 94.41
$\text{CNN}_{\text{large}}$ 6.67 0.04 0.05 0.04 14.08 8.23 7.83 92.72
Conversely, for the simple dataset (MNIST), CKANs exhibit significantly better robustness than CNNs across all model sizes. Notably, medium and large CKAN models demonstrate remarkable resilience, successfully classifying a significant proportion of adversarial images. In contrast, CNNs perform poorly under these conditions, managing to classify only a negligible number of adversarial images.
This disparity is further corroborated by Figure 3, which reveals a faster loss convergence on complex datasets (CIFAR-10 and SVHN) compared to MNIST during the attack process. The rapid convergence of the loss implies that adversarial attacks “fool” models more easily on these complex datasets.
Figure 3: Comparative loss dynamics of CNNs and CKANs under PGD, MIM, and C&W attacks on the MNIST, CIFAR-10, and SVHM datasets. Each line represents the mean loss across batches per iteration, with shaded areas indicating the standard deviation.
For the non-iterative white-box attack (FGSM), the results are more definitive. On CIFAR-10, CKANs exhibit greater robustness than CNNs for medium-sized models, whereas CNNs perform better for large models. On the SVHN dataset, CKANs demonstrate slightly higher robustness across all model sizes, although the gap between CKANs and CNNs is relatively small. On the MNIST dataset, which is less complex, CKANs show significantly higher robustness than CNNs, with a noticeable gap favoring CKANs. As shown in Figure 4, this gap is not consistently large across different $\epsilon$ values.
Figure 4: Robust accuracy of $\text{CNN}_{\text{medium}}$, $\text{CNN}_{\text{large}}$, $\text{CKAN}_{\text{medium}}$, and $\text{CKAN}_{\text{large}}$ as a function of FGSM attack strength, with varying $\epsilon$ values. Each bar represents a different model, and robustness is measured as the accuracy of the model against adversarial examples generated with specific $\epsilon$ values. The x-axis indicates the $\epsilon$ values used for FGSM attacks, while the y-axis shows the corresponding robust accuracy of the models.
When analyzing black-box attacks, CKANs consistently demonstrated superior robustness compared to CNNs across most datasets and model sizes. CKANs outperformed CNNs in nearly all scenarios, except for the black-box attacks on large models for the CIFAR-10 dataset, where CNNs exhibited slightly better robustness, and on the small models for the CIFAR-10 dataset, where both of the configurations failed to classify any adversarial image generated with the Square attack.
In our analysis of the transferability of adversarial white-box attacks (Table 7), we observed that the results vary across datasets and model configurations, with no clear overall trend. On the MNIST dataset, fooling the medium CKAN proves to be the most challenging task. On the SVHN dataset, the large CKAN is the hardest modle to fool. Conversely, on the CIFAR-10 dataset, the large CKAN is the easiest model to fool, right after the medium CNN model. Similar to the transferability results for the fully connected models, CKANs consistently demonstrated greater resistance to adversarial attacks than their CNN counterparts for each model size, with the exception of the large size on the CIFAR-10 dataset.
Table 7: Maximum transferability of our convolutional models for MNIST, CIFAR-10, and SVHM. The value in row $i$, column $j$ represents the maximum transferability between row-$i$ model and column-$j$ model across four attacks—FGSM, PGD, C&W, MIM—calculated using Equation 11. Unlike robustness tables, the values here indicate the percentage of images misclassified. The ‘Average’ row shows the average transferability for the attacking model. Boldface values denoting the lowest transferability, where a lower value indicates better robustness.
(a) MNIST
$\text{CKAN}_{\text{small}}$ $\text{CKAN}_{\text{medium}}$ $\text{CKAN}_{\text{large}}$ $\text{CNN}_{\text{small}}$ $\text{CNN}_{\text{medium}}$ $\text{CNN}_{\text{large}}$
$\text{CKAN}_{\text{small}}$ 95.21 2.94 3.26 4.70 20.6 4.80
$\text{CKAN}_{\text{medium}}$ 5.30 74.81 17.78 3.32 18.95 12.01
$\text{CKAN}_{\text{large}}$ 3.75 14.95 72.80 2.51 15.16 9.93
$\text{CNN}_{\text{small}}$ 3.43 1.71 1.55 99.22 33.47 6.72
$\text{CNN}_{\text{medium}}$ 3.82 2.78 3.48 7.80 100.00 46.28
$\text{CNN}_{\text{large}}$ 3.35 4.49 5.28 5.64 6.98 100.00
Average ($\downarrow$) 19.14 16.94 17.35 20.53 32.52 29.95
(b) CIFAR-10
$\text{CKAN}_{\text{small}}$ $\text{CKAN}_{\text{medium}}$ $\text{CKAN}_{\text{large}}$ $\text{CNN}_{\text{small}}$ $\text{CNN}_{\text{medium}}$ $\text{CNN}_{\text{large}}$
$\text{CKAN}_{\text{small}}$ 100.00 53.39 45.62 62.95 60.19 35.45
$\text{CKAN}_{\text{medium}}$ 70.25 100.00 93.48 79.12 90.40 68.52
$\text{CKAN}_{\text{large}}$ 62.80 92.13 100.00 75.85 89.65 73.39
$\text{CNN}_{\text{small}}$ 70.47 71.27 69.74 100.00 85.70 59.78
$\text{CNN}_{\text{medium}}$ 63.23 82.31 82.52 82.50 100.00 78.68
$\text{CNN}_{\text{large}}$ 47.53 68.93 76.86 65.07 84.74 100.00
Average ($\downarrow$) 69.04 78.00 78.03 77.58 85.15 69.30
(c) SVHN
$\text{CKAN}_{\text{small}}$ $\text{CKAN}_{\text{medium}}$ $\text{CKAN}_{\text{large}}$ $\text{CNN}_{\text{small}}$ $\text{CNN}_{\text{medium}}$ $\text{CNN}_{\text{large}}$
$\text{CKAN}_{\text{small}}$ 100.00 63.08 58.62 72.23 58.73 57.70
$\text{CKAN}_{\text{medium}}$ 77.44 100.00 83.76 69.97 75.40 75.40
$\text{CKAN}_{\text{large}}$ 76.54 86.43 99.90 69.42 76.23 75.67
$\text{CNN}_{\text{small}}$ 68.52 55.79 52.85 100.00 63.33 62.15
$\text{CNN}_{\text{medium}}$ 72.11 74.21 71.98 78.26 99.96 85.87
$\text{CNN}_{\text{large}}$ 70.93 74.52 72.70 78.62 85.89 99.96
Average ($\downarrow$) 77.59 75.67 73.30 78.08 76.58 76.12
5 Concluding Remarks
The introduction of KANs opens new avenues in adversarial machine learning, presenting both challenges and opportunities for improving model security. In this study, we explored several key aspects of KANs, focusing on their behavior in convolutional and fully connected layers, which have not been extensively investigated before.
First, we assessed the inherent robustness of FCKANs and CKANs against both white-box and black-box adversarial attacks. Our findings indicate that while FCKANs exhibit vulnerability levels similar to traditional FCNNs in smaller configurations, they demonstrate significant improvements in robustness as network size increases. This suggests that FCKANs could be particularly effective in scenarios where larger models are feasible and robustness is a priority. In contrast to the trends observed for FCKANs and FCNNs, CKANs consistently exhibit greater robustness than CNNs of comparable sizes across most attacks. This highlights a distinct advantage of CKANs in adversarial settings, even when network sizes remain constant.
Despite these promising findings, it is crucial to recognize that KANs remain vulnerable to adversarial perturbations. Our analysis reveals that while larger configurations enhance robustness, KANs are not impervious to adversarial attacks. This ongoing fragility underscores the need for continued research to develop more effective strategies for improving the security of KANs. Addressing these vulnerabilities is essential for the reliable deployment of KANs in security-critical applications.
As the field continues to advance, further investigation is required to better understand the strengths and limitations of FCKANs and CKANs. This research will be instrumental in paving the way for the practical application of KANs in domains where security and reliability are paramount.
Appendix A Image Samples of Attacks
This section provides visual examples of adversarial attacks applied to a single MNIST image, demonstrating the unique perturbation patterns across different model architectures and sizes. These visualizations compare FCNNs with FCKANs and CNNs with CKANs under white-box and black-box attack scenarios.
Fully Connected Models. Figure 5 and Figure 6 depict adversarial images and their corresponding perturbations for FCNNs and FCKANs under white-box and black-box attacks, respectively. The results highlight distinct perturbation behaviors, underscoring the unique adversarial behaviors of KANs.
Convolutional Models. Figure 7 and Figure 8 illustrate adversarial images and perturbations for CNNs and CKANs under white-box and black-box attacks, showcasing architectural influences on adversarial patterns.
Figure 5: Visualization of the adversarial images on a single MNIST image using fully connected models. Rows represent model types and sizes: $\text{FCNN}_{\text{large}}$, $\text{FCKAN}_{\text{large}}$, $\text{FCNN}_{\text{medium}}$, $\text{FCKAN}_{\text{medium}}$, $\text{FCNN}_{\text{small}}$, and $\text{FCKAN}_{\text{small}}$.
Figure 6: Visualization of the pertubations on a single MNIST image using fully connected models. Rows represent model types and sizes: $\text{FCNN}_{\text{large}}$, $\text{FCKAN}_{\text{large}}$, $\text{FCNN}_{\text{medium}}$, $\text{FCKAN}_{\text{medium}}$, $\text{FCNN}_{\text{small}}$, and $\text{FCKAN}_{\text{small}}$.
Figure 7: Visualization of the adversarial images on a single MNIST image using convolutional models. Rows represent model types and sizes: $\text{CNN}_{\text{large}}$, $\text{CKAN}_{\text{large}}$, $\text{CNN}_{\text{medium}}$, $\text{CKAN}_{\text{medium}}$, $\text{CNN}_{\text{small}}$, and $\text{CKAN}_{\text{small}}$.
Figure 8: Visualization of the pertubations on a single MNIST image using MNIST image using convolutional models. Rows represent model types and sizes: $\text{CNN}_{\text{large}}$, $\text{CKAN}_{\text{large}}$, $\text{CNN}_{\text{medium}}$, $\text{CKAN}_{\text{medium}}$, $\text{CNN}_{\text{small}}$, and $\text{CKAN}_{\text{small}}$.
Appendix B Adversarial Attacks
In this appendix we provide a formal mathematical definition of the white-box and black box adversarial attacks.
B.1 White-Box Attacks
B.1.1 Fast Gradient Sign Method (FGSM)
The Fast Gradient Sign Method (FGSM) (Goodfellow et al., 2014) generates adversarial examples by adding perturbations in the direction of the gradient of the loss function. Specifically, an adversarial example $x_{\text{adv}}$ is computed as follows:
where $\mathcal{L}$ is the loss function associated with a classifier parameterized by $\theta$, $\nabla_{x}\mathcal{L}(x,y;\theta)$ represents the gradient of the loss with respect to the input $x$, and $\epsilon$ is a small perturbation. Notably, in Equation 1, only the sign of the gradient is used, with the magnitude of the perturbation controlled by $\epsilon$. It is crucial to highlight that FGSM is a single-step attack; the adversary computes the gradient by backpropagating through the model once—and directly applies this gradient to perturb the input $x$.
B.1.2 Projected Gradient Descent (PGD)
The Projected Gradient Descent (PGD) method (Madry et al., 2017) is an iterative extension of the FGSM, designed to generate more robust adversarial examples. The PGD attack iteratively perturbs the input $x$ by repeatedly applying gradient updates and projecting the perturbed image back onto an $\epsilon$-ball around the original input. The adversarial example $x_{\text{adv}}$ after $k$ iterations is computed as:
B.1.3 Carlini & Wagner (C&W)
The Carlini & Wagner (C&W) is a powerful optimization-based adversarial attack designed to generate adversarial examples by minimizing a specially crafted loss function while ensuring that the perturbation remains imperceptible. When employing the $\ell_{\infty}$ norm, the attack seeks to find a minimal perturbation $\delta$ such that the resulting adversarial example $x_{\text{adv}}=x+\delta$ misleads the classifier. The optimization problem is formulated as follows:
$\mathcal{L}(\delta,x)=\|\delta\|_{\infty}+c\cdot f(x+\delta),$ (3)
where $\mathcal{L}(\delta,x)$ is the objective function to minimize, subject to:
$x_{\text{adv}}=\text{clip}(x+\delta,0,1).$ (4)
The attack is typically solved through an iterative optimization procedure, where the perturbation $\delta$ and the adversarial image $x_{\text{adv}}$ are updated at each iteration, as follows:
$\delta^{k}$ $=\text{clip}(x_{\text{adv}}^{k}-x,-\epsilon,\epsilon),$ (5)
$x_{\text{adv}}^{k+1}$ $=x_{\text{adv}}^{k}+\alpha\cdot\text{sign}\bigl{(}\nabla_{x}\mathcal{L}(x_{\text{adv}}^{k},\delta^{k})\bigr{)}.$
where $\alpha$ is the step size, and $k$ indicates the iteration number.
B.1.4 Momentum Iterative Method (MIM)
The Momentum Iterative Method (MIM) (Dong et al., 2018) is an enhancement of the basic iterative methods for generating adversarial examples, incorporating momentum to stabilize the gradient updates and avoid poor local maxima. The MIM attack iteratively updates the adversarial example $x_{\text{adv}}$ by accumulating a momentum term, which helps to amplify gradients that consistently point in the same direction across iterations. The adversarial example $x_{\text{adv}}^{k+1}$ after $k$ iterations is computed as:
B.2 Black-Box Attacks
In a black-box attack the adversary does not have direct access to the internal parameters or architecture of the target model. Instead, the attack relies on querying the model to gather information about its predictions or confidence scores. Adversarial examples $x_{\text{adv}}$ are crafted based on these outputs, without explicit knowledge of the model’s gradients or weights. Below, we describe the specific black-box attack methodologies used in this study.
B.2.1 Transfer-based Attacks
Transfer-based attacks leverage the phenomenon of transferability, where adversarial examples generated for one model are also misclassified by other models. This concept of adversarial example transferability was first introduced by Szegedy et al. (2013).